# Minimal Generation of Mapping Class Groups: A Survey of the Orientable Case

### v1

@@ 2.2. Dehn Twists: Fundamental Generators
$t_{a}(x)=\begin{cases}h^{-1}\circ t\circ h(x)&x\in N\\ x&x\in\Sigma_{g}\setminus N\end{cases}$
@@ 2.4. Elementary Mapping Class Groups
In the study of surfaces and their symmetries, two other fundamental objects emerge: the Teichmüller space and the moduli space. These interconnected concepts play crucial roles in geometric group theory, particularly in the investigation of generating sets for mapping class groups. Before we finish this section let us give their definitions, relationships, and importance in the context of geometric group theory. Let $\Sigma$ denote an orientable surface that is either closed, have punctures, or have boundary components.
@@ 4. GENERATION BY TORSION ELEMENTS
    Lanier [14] proved that for any integer $k\geq 6$, the mapping class group $\operatorname{Mod}(\Sigma_{g})$ is generated by three elements of order $k$ provided that the genus $g$ satisfies $g\geq 2((k-1)^{2}+1)$. His construction begins with a $k$-fold symmetric embedding $S_{g}$ of $\Sigma_{g}$ in $\mathbb{R}^{3}$ where a rotation $r$ of $\Sigma_{g}$ by an angle of $2\pi/k$ about its axis yields an element of order $k$ similar to [22]. To obtain the remaining elements of order $k$, Lanier embeds $k$-symmetric subsurfaces $\sigma_{i}$ into $\Sigma_{g}$ into and within each $\sigma_{i}$, he places a chain of simple closed curves of length $2g_{i}-1$, where $g_{i}$ is the genus of $\sigma_{i}$. He then constructs the required elements by mapping $S_{g}$ to $\Sigma_{g}$, applying the rotation $r$, and mapping back. This careful use of symmetry and surface decomposition yields three torsion elements of order $k$ that generate the entire group.
@@ 4.1. Involution Generators
The proof that the mapping class group $\operatorname{Mod}_{g}$ is generated by $6$ involutions is constructive. The strategy is to start with a known, small set of generators and show that each can be expressed as a product of elements from a set of just six involutions. The authors begin with a known generating set for $\operatorname{Mod}_{g}$ consisting of four elements: three Dehn twists, $t_{\alpha}$, $t_{\beta}$, $t_{\gamma}$, and a rotation $R_{g}$. The rotation $R_{g}$ together with the three twists generate all of Lickorish’s 3g13g-1 twist generators. The core of the proof is to express these four generators using a small, common set of involutions.
@@
[10] Let $\Sigma_{g}$ be a closed, oriented surface of genus $g\geq 7$. There exist involutions
@@ 5.1. The Birman Exact Sequence and Point-Pushing Maps
    Start with known generators for $\mathrm{PMod}(\Sigma_{g,0})=\operatorname{Mod}{\Sigma_{g}}$ (if $i=0$) or $\mathrm{PMod}(\Sigma_{0,p})$ (if $g=0$).
@@ 5.3.2. The Case of Multiple Punctures
These results are significant as they provide concrete numbers of involution generators for mapping class groups of punctured surfaces, contributing to the broader understanding of their algebraic structure and the result in the first case is sharp since the mapping class group is not generated by two involutions. The proofs of these statements often involve intricate geometric constructions of the involutions and detailed analysis of their products, showing that they can be combined to produce a known set of generators for the mapping class group, such as a set of Dehn twists. During the writing of this chapter, it was found that Theorem 5.8 (i) holds true even when $g=13$, representing a meaningful improvement over our prior results. The proof that follows not only establishes this improvement but also lays out the same method used to prove the original Theorem 5.8. We consider embeddings of the surface $\Sigma_{g,p}$ in $\mathbb{R}^{3}$ that remain invariant under the $\pi$ rotations $\rho_{1}$ and $\rho_{2}$ about the z-axis (shown in [1, Figure 1] for $g=2k$ and Figure 10 for $g=2k+1$). The mapping class group $\operatorname{Mod}(\Sigma_{g,p})$ contains the element $R=\rho_{1}\rho_{2}$ with the following properties:
@@ 6.1. Mapping Class Groups of Simple Surfaces with Boundary
$\operatorname{Mod}(\Sigma_{0}^{1})=\{\operatorname{id}\}$
@@
$\operatorname{Mod}(\Sigma_{0}^{2})\cong\mathbb{Z}$
@@
Let $P=\Sigma_{0}^{3}$ be a pair of pants (a sphere with three disjoint open disks removed). If all three boundary components are fixed pointwise, the mapping class group is trivial:
$\operatorname{Mod}(\Sigma_{0}^{3})=\{\operatorname{id}\}.$
Any essential simple closed curve in the interior of a pair of pants is parallel to one of the boundary components. A Dehn twist about such a curve, while keeping the boundary fixed, would necessarily be supported in an annulus bounded by this curve and the corresponding boundary component. Such a twist is isotopic to the identity if the boundary itself is fixed pointwise. The situation is analogous to twisting an annulus while holding one of its boundary circles perfectly still, an action which can always be unwound.
@@ 6.2. Generators for General Surfaces with Boundary
[6] The mapping class group $\operatorname{Mod}(\Sigma^{1}_{g})$ of an orientable surface of genus $g$ with one boundary component is generated by $2g$ Dehn twists.
@@ 6.3. Relationship with Mapping Class Groups of Closed Surfaces
    Capping with a Disk: If we glue a disk $D$ to a boundary component $\partial_{i}S$ of $\Sigma^{b}_{g}$ to obtain $\Sigma_{g}^{b-1}$ (or $\Sigma_{g}$ if $b=1$), there is a surjective homomorphism: cap_i : Mod(Σ^b_g) →Mod(Σ_g^b-1) The kernel of this map is generated by the Dehn twist $t_{c_{i}}$ about a curve $c_{i}$ parallel to the boundary component $\partial_{i}S$ that was capped. This leads to a short exact sequence (for $b\geq 1$ and assuming $\Sigma_{g}^{b-1}$ is not $\Sigma^{0}_{0}$ or $\Sigma^{1}_{0}$ which have trivial $\operatorname{Mod}$): 1 →⟨t_c_i ⟩≅Z →ModΣ(g, b) →Mod(Σ_g, b-1) →1. This sequence can be used inductively. Repeatedly capping all $b$ boundary components relates $\operatorname{Mod}(\Sigma_{g}^{b})$ to $\operatorname{Mod}(\Sigma_{g})$. The kernel in this case would be more complex, involving multiple boundary twists.
@@
    Forgetting Punctures (if $n>0$): The Birman exact sequence relates $\operatorname{Mod}(\Sigma_{g,n}^{b})$ to $\operatorname{Mod}(\Sigma_{g,n-1}^{b})$: 1 →π_1(Σ^b_g, n-1, p_n) push→ Mod(Σ_g, n^b) →Mod(Σ_g, n-1^b) →1. The generators for the kernel are point-pushing maps, expressible as Dehn twists.

### v2

@@ 2.2. Dehn Twists: Fundamental Generators
$t_{a}(x)=\begin{cases}h^{-1}\circ t\circ h(x)&x\in N\\ x&x\in\Sigma_{g}\setminus N,\end{cases}$
@@ 2.4. Elementary Mapping Class Groups
In the study of surfaces and their symmetries, two other fundamental objects emerge: the Teichmüller space and the moduli space. These interconnected concepts play crucial roles in geometric group theory, particularly in the investigation of generating sets for mapping class groups. Before we finish this section, let us give their definitions, relationships, and importance in the context of geometric group theory. Let $\Sigma$ denote an orientable surface that is either closed, has punctures, or has boundary components.
@@ 4. GENERATION BY TORSION ELEMENTS
    Lanier [14] proved that for any integer $k\geq 6$, the mapping class group $\operatorname{Mod}(\Sigma_{g})$ is generated by three elements of order $k$ provided that the genus $g$ satisfies $g\geq 2((k-1)^{2}+1)$. His construction begins with a $k$-fold symmetric embedding $S_{g}$ of $\Sigma_{g}$ in $\mathbb{R}^{3}$ where a rotation $r$ of $\Sigma_{g}$ by an angle of $2\pi/k$ about its axis yields an element of order $k$ similar to [22]. To obtain the remaining elements of order $k$, Lanier embeds $k$-symmetric subsurfaces $\sigma_{i}$ into $\Sigma_{g}$ and within each $\sigma_{i}$, he places a chain of simple closed curves of length $2g_{i}-1$, where $g_{i}$ is the genus of $\sigma_{i}$. He then constructs the required elements by mapping $S_{g}$ to $\Sigma_{g}$, applying the rotation $r$, and mapping back. This careful use of symmetry and surface decomposition yields three torsion elements of order $k$ that generate the entire group.
@@ 4.1. Involution Generators
The proof that the mapping class group $\operatorname{Mod}(\Sigma_{g})$is generated by $6$ involutions is constructive. The strategy is to start with a known, small set of generators and show that each can be expressed as a product of elements from a set of just six involutions. The authors begin with a known generating set for $\operatorname{Mod}(\Sigma_{g})$ consisting of four elements: three Dehn twists, $t_{\alpha}$, $t_{\beta}$, $t_{\gamma}$, and a rotation $R_{g}$. The rotation $R_{g}$ together with the three twists generate all of Lickorish’s 3g13g-1 twist generators. The core of the proof is to express these four generators using a small, common set of involutions.
@@
[10] Let $\Sigma_{g}$ be a closed, oriented surface of genus $g\geq 7$. There exist elements
@@ 5.1. The Birman Exact Sequence and Point-Pushing Maps
    Start with known generators for $\mathrm{PMod}(\Sigma_{g,0})=\operatorname{Mod}(\Sigma_{g})$ (if $i=0$) or $\mathrm{PMod}(\Sigma_{0,p})$ (if $g=0$).
@@ 5.3.2. The Case of Multiple Punctures
These results are significant as they provide concrete numbers of involution generators for mapping class groups of punctured surfaces, contributing to the broader understanding of their algebraic structure and the result in the first case is sharp since the mapping class group is not generated by two involutions. The proofs of these statements often involve intricate geometric constructions of the involutions and detailed analysis of their products, showing that they can be combined to produce a known set of generators for the mapping class group, such as a set of Dehn twists. In this chapter, we improve upon our prior results by demonstrating that Theorem 5.8 (i) also holds for $g=13$. The proof that follows not only establishes this improvement but also lays out the same method used to prove the original Theorem 5.8. We consider embeddings of the surface $\Sigma_{g,p}$ in $\mathbb{R}^{3}$ that remain invariant under the $\pi$ rotations $\rho_{1}$ and $\rho_{2}$ about the z-axis (shown in [1, Figure 1] for $g=2k$ and Figure 10 for $g=2k+1$). The mapping class group $\operatorname{Mod}(\Sigma_{g,p})$ contains the element $R=\rho_{1}\rho_{2}$ with the following properties:
@@ 6.1. Mapping Class Groups of Simple Surfaces with Boundary
$\operatorname{Mod}(\Sigma_{0}^{1})=\{\operatorname{id}\}.$
@@
$\operatorname{Mod}(A)\cong\mathbb{Z}.$
@@
Let $P=\Sigma_{0}^{3}$ be a pair of pants (a sphere with three disjoint open disks removed). The mapping class group $\operatorname{Mod}(P)$ is nontrivial:
$\operatorname{Mod}(P)\cong\mathbb{Z}^{3}.$
This group is generated by Dehn twists about the three boundary-parallel curves which clearly commute. Hence there is an injective homomorphism $\mathbb{Z}^{3}\to\operatorname{Mod}(P)$. By applying the Birman exact sequence to each boundary circle yields a surjection $\operatorname{Mod}(P)\to\operatorname{Mod}(A)\cong\mathbb{Z}$ whose kernel is generated by the Dehn twist about the capped boundary, we see that the subgroup generated by the three boundary-parallel Dehn twists equals $\operatorname{Mod}(P)$.
@@ 6.2. Generators for General Surfaces with Boundary
[9] The mapping class group $\operatorname{Mod}(\Sigma^{1}_{g})$ of an orientable surface of genus $g$ with one boundary component is generated by $2g+1$ Dehn twists.
@@ 6.3. Relationship with Mapping Class Groups of Closed Surfaces
    Capping with a Disk: If we glue a disk $D$ to a boundary component $\partial_{i}S$ of $\Sigma^{b}_{g}$ to obtain $\Sigma_{g}^{b-1}$ (or $\Sigma_{g}$ if $b=1$), there is a surjective homomorphism:
    $\text{cap}_{i}:\operatorname{Mod}(\Sigma^{b}_{g})\to\operatorname{Mod}(\Sigma_{g}^{b-1})$
    The kernel of this map is generated by the Dehn twist $t_{c_{i}}$ about a curve $c_{i}$ parallel to the boundary component $\partial_{i}S$ that was capped. This leads to a short exact sequence (for $b\geq 1$ and assuming $\Sigma_{g}^{b-1}$ is not $\Sigma^{0}_{0}$ or $\Sigma^{1}_{0}$ which have trivial $\operatorname{Mod}$):
    $1\to\langle t_{c_{i}}\rangle\cong\mathbb{Z}\to\operatorname{Mod}\Sigma({g,b})\to\operatorname{Mod}(\Sigma_{g,b-1})\to 1.$
    This sequence can be used inductively. Repeatedly capping all $b$ boundary components relates $\operatorname{Mod}(\Sigma_{g}^{b})$ to $\operatorname{Mod}(\Sigma_{g})$. The kernel in this case would be more complex, involving multiple boundary twists.
@@
    Forgetting Punctures (if $n>0$): The Birman exact sequence relates $\operatorname{Mod}(\Sigma_{g,n}^{b})$ to $\operatorname{Mod}(\Sigma_{g,n-1}^{b})$:
    $1\to\pi_{1}(\Sigma^{b}_{g,n-1},p_{n})\xrightarrow{\text{push}}\operatorname{Mod}(\Sigma_{g,n}^{b})\to\operatorname{Mod}(\Sigma_{g,n-1}^{b})\to 1.$
    The generators for the kernel are point-pushing maps, expressible as Dehn twists.
@@ 7.1. Future Directions
Acknowledgements The authors thank Athanase Papadopoulos for inviting them to write this survey and Allen Hatcher for correcting a mistake in an earlier version of this work.